# Imputation-free Learning of Tabular Data with Missing Values using Incremental Feature Partitions in Transformer

### v4

@@ 4 Results
All results are obtained using an Ubuntu 22.04 machine utilizing an Intel(R) Xeon(R) W-2265 CPU (24 logical cores) running at 3.70GHz, 64GB of RAM, and a Quadro RTX A4000 GPU with 16GB of video memory. The performance of the proposed NIIL method is compared with the baseline methods using 1) the average rank based on AUC scores, 2) the win matrix for pairwise method comparison, and 3) the robustness of the performance to increasing missing rates. B presents the AUC scores specific to the 13 data sets of type MCAR in Table 6 and of type MNAR in Table 7.
@@
Table 2: Average performance rank of the proposed (NIIL) and baseline models across 13 diverse MCAR-type data sets with five missing value rates.
Method Imputation 10% 20% 30% 40% 50% Avg. Rank Overall
Median-GBT Yes 7.8(0.05) 7.4(0.05) 7.5(0.04) 8.3(0.06) 7.2(0.05) 7.7(0.05) 7
Median-FTT Yes 8.7(0.07) 8.7(0.07) 8.9(0.06) 9.0(0.06) 9.8(0.07) 9.0(0.07) 12
MICE-GBT Yes 6.5(0.05) 6.8(0.05) 8.3(0.05) 8.2(0.05) 7.8(0.06) 7.5(0.05) 6
MICE-FTT Yes 8.0(0.05) 7.6(0.05) 6.9(0.04) 6.8(0.04) 7.2(0.05) 7.3(0.05) 5
GAIN-GBT Yes 10.0(0.05) 9.2(0.05) 7.4(0.05) 7.1(0.06) 6.6(0.04) 8.0(0.05) 8
GAIN-FTT Yes 7.8(0.06) 6.6(0.06) 6.8(0.06) 7.1(0.07) 6.8(0.06) 7.0(0.06) 4
Diffputer-GBT Yes 9.1(0.06) 9.1(0.06) 10.2(0.06) 11.9(0.06) 9.7(0.06) 10.0(0.06) 14
Diffputer-FTT Yes 8.7(0.06) 9.7(0.06) 10.0(0.06) 9.3(0.06) 10.1(0.06) 9.5(0.06) 13
MIA-Xgboost No 9.1(0.05) 8.9(0.05) 8.4(0.05) 8.9(0.05) 8.2(0.05) 8.7(0.05) 11
MIA-LightGBM No 8.5(0.04) 9.2(0.04) 9.0(0.05) 9.5(0.05) 7.2(0.05) 8.7(0.05) 10
AM-FTT No 9.3(0.06) 10.5(0.06) 10.8(0.06) 10.3(0.07) 11.5(0.07) 10.5(0.06) 15
NIIL (K =2) No 8.2(0.04) 8.2(0.05) 7.6(0.04) 7.8(0.05) 8.5(0.06) 8.0(0.05) 9
NIIL (K =3) No 6.7(0.05) 6.0(0.04) 5.9(0.04) 5.9(0.04) 6.2(0.05) 6.1(0.05) 2
NIIL (K =4) No 6.4(0.04) 6.0(0.04) 6.5(0.05) 5.2(0.04) 7.0(0.05) 6.2(0.05) 3
NIIL (K =d/2) No 5.5(0.05) 6.1(0.05) 5.8(0.05) 4.8(0.04) 6.2(0.05) 5.7(0.05) 1
@@
Table 3: Average performance rank of the proposed (NIIL) and baseline models across 13 diverse MNAR-type data sets with five missing value rates.
Method Imputation 10% 20% 30% 40% 50% Avg. Rank Overall
Median-GBT Yes 7.2(0.05) 7.3(0.06) 7.3(0.06) 8.5(0.05) 5.8(0.06) 7.2(0.06) 6
Median-FTT Yes 7.4(0.05) 7.2(0.05) 6.5(0.06) 7.7(0.06) 8.8(0.06) 7.5(0.06) 9
MICE-GBT Yes 9.1(0.05) 8.0(0.04) 7.0(0.06) 7.2(0.06) 5.8(0.05) 7.4(0.05) 8
MICE-FTT Yes 11.1(0.05) 9.5(0.06) 10.1(0.06) 9.5(0.06) 10.0(0.07) 10.0(0.06) 13
GAIN-GBT Yes 8.5(0.04) 8.8(0.05) 9.9(0.04) 8.6(0.04) 8.8(0.04) 8.9(0.04) 12
GAIN-FTT Yes 11.5(0.07) 11.9(0.08) 11.2(0.06) 12.3(0.06) 12.9(0.06) 12.0(0.06) 15
Diffputer-GBT Yes 9.2(0.06) 10.5(0.07) 10.9(0.06) 10.0(0.05) 9.8(0.06) 10.1(0.06) 14
Diffputer-FTT Yes 8.4(0.07) 7.5(0.06) 8.8(0.07) 8.0(0.06) 10.1(0.06) 8.6(0.06) 11
MIA-Xgboost No 5.5(0.05) 6.8(0.05) 7.5(0.05) 8.5(0.05) 7.2(0.06) 7.1(0.05) 5
MIA-LightGBM No 6.5(0.05) 7.7(0.04) 7.2(0.05) 7.0(0.05) 7.9(0.06) 7.3(0.05) 7
AM-FTT No 9.9(0.06) 7.5(0.06) 8.2(0.05) 8.9(0.05) 8.2(0.06) 8.6(0.06) 10
NIIL (K =2) No 7.2(0.05) 7.1(0.06) 6.3(0.06) 6.4(0.05) 5.7(0.06) 6.5(0.06) 4
NIIL (K =3) No 6.6(0.05) 6.2(0.06) 6.8(0.07) 5.9(0.05) 6.2(0.06) 6.3(0.06) 2
NIIL (K =4) No 6.2(0.05) 7.2(0.06) 6.3(0.06) 5.7(0.05) 6.4(0.06) 6.4(0.06) 3
NIIL (K =d/2) No 5.8(0.04) 6.7(0.07) 6.1(0.06) 5.7(0.06) 6.4(0.06) 6.1(0.06) 1
@@
Table 4: Mean AUC scores after classifying data sets (Hepatitis (ID: 55), Cylinder-bands (ID: 6332)) with naturally occurring missing values at the rate of 48.8% and 48.7%, respectively.
ID Median- Median- MICE- MICE- GAIN- GAIN- Diffputer- Diffputer- MIA- MIA- AM-FTT NIIL
@@ 4.1 Average performance rank
Table 2 compares the proposed NIIL method with 11 baselines for MCAR-type data. In single-value imputation, median imputation is better when GBT is a classifier than FTT, regardless of missing rates. The MICE or multiple-imputation-based method outperforms median value imputation. The GAN-based imputation method (GAIN) with FTT performs better than MICE for all missing rates. In contrast, the diffusion-based imputation method performs the worst among all imputation techniques. Among the methods without explicit imputations, the MIA-based models (MIA-XGBoost and MIA-LightGBM) perform better than AM-FTT. The average performance rank shows that the proposed NIIL (K = d/2) outperforms all 11 baselines. NIIL with other feature partition sizes, including NIIL (K = 3) and NIIL (K = 4), also rank among the top three. Even at low missing rates (10%–30%), NIIL (K = d/2) maintains a strong performance rank, while at higher rates it demonstrates clear superiority. These results confirm the robustness and consistent performance of the proposed NIIL method for MCAR-type data.
Table 3 presents the average classification performance ranks in classifying MNAR-type data. Imputation-based methods that present competitive performance for MCAR-type data (GAIN-FTT and MICE-FTT) show a substantial performance drop for MNAR-type data, ranking 15th and 13th, respectively. In contrast, MIA-based models (MIA-XGBoost and MIA-LightGBM) perform better on MNAR data than MCAR data. This observation is consistent with previous findings [41] suggesting the MIA method performs better when the missingness is informative, such as MNAR. At lower missing rates (10% and 20%), the average ranks suggest that MIA methods outperform all methods that require an imputation step. However, median-GBT consistently outperforms MIA and other imputation-based methods when the missing rate is above 20%. On the other hand, NIIL with other partition sizes (K = 2, 3 and 4) is among the top four best methods, while (K = d / 2) is the best performing method. These results suggest that the proposed NIIL method is not impacted by missing value types, unlike other imputation-based approaches. Table 4 presents the performance of two data sets with naturally occurring missing values, Hepatitis (ID: 55) and Cylinder-bands (ID: 6332) with missing rates of 48.8% and 48.7%, respectively. The proposed NIIL method (K = d/2) achieves the best average performance rank (2.5 (0.03)), while the second best method is median-GBT (average rank 4.0 (0.05). This again shows the robustness of the NIIL method to missing value types.
@@
Figure 2: Win matrix. Values are the fraction of experimental scenarios in which the row methods outperform methods in the columns.
4.2 Pairwise model comparison
We count the experimental scenarios in which one method wins over another. There are 65 experimental scenarios from 13 data sets and five missing rates.
Figure 2 (a) for the MCAR-type data shows that the proposed NIIL method is always superior to the baselines for more than 50% of the experimental scenarios. In the best case, NIIL outperforms Diffputer-FTT and MIA-Xgboost in 72% of the scenarios. Figure 2 (b) for the MNAR-type data shows that NIIL is always superior to the baselines for more than 50% of the experimental scenarios. In the best case, NIIL outperforms GAIN-FTT in 83% of the scenarios, which is 52% for MCAR-type data. This indicates that deep-imputation methods are not optimized for MNAR-type data.
@@ 4.3 Computational time and robustness to missing rates
Table 5 presents the run time for the imputation and classification steps of the baseline methods. As expected, methods faster than NIIL include GAIN-GBT, Diffputer-GBT, AM-FTT, MIA-Xgboost, and MIA-LightGBM, mainly due to the fast GBT classifier model. However, the classification performance rank of these methods is worse than that of NIIL. One computational benefit of NIIL is that it saves the time required for imputation. Overall, the run-time is decent for tabular data sets of standard sample sizes (2̃000).
Figure 3 shows the effect of varying the missing rates on the performance of five top-performing models. The reference AUC scores for the robustness analysis are obtained from complete data sets with 0% missing values. The AUC score drops against the reference AUC due to increased missing rates. Figure 3(a) for MCAR-type data shows the robustness of the proposed method when the missing rate is higher than 30%. GAIN-FTT is the most robust method up to 30% missing values. However, Figure 3(b) for MNAR-type data shows that the robustness of the proposed NIIL method to high missing rates is better than all other baselines.
@@
MICE - GBT 417 2 419 6 7
MICE - FTT 417 190 607 5 13
GAIN - GBT 51 2 54 8 11
GAIN - FTT 51 190 241 3 15
Diffputer - GBT 8 1 9 13 14
Diffputer - FTT 8 190 198 12 10
AM-FTT N/A 21 21 15 12
MIA - Xgboost N/A 1 1 10 6
MIA - LightGBM N/A 1 1 9 8
NIIL (k = $\frac{d}{2}$) N/A 163 163 1 1
@@ 4.4 Ablation studies
In an ablation study, we vary the value of feature partition size k for incremental learning using three representative data sets: Kc2 (ID 1063), Diabetes (ID 37), and Dresses sales (ID 23381). Figure 4 shows the k values that yield the best AUC scores for individual data sets. In general, feature incremental learning is superior to learning with all features at a time (i.e., no incremental learning, k = d). The best average AUCs are obtained for the data sets Kc2 $(d=21)$ at k=10, Diabetes $(d=8)$ at k=6, and Dresses sales $(d=12)$ at k = 6. Therefore, the value of k can be chosen as half of the feature dimension for the best classification performance. This observation aligns with the computational cost analysis presented in Figure 5 in  A, where $k<\frac{d}{2}$ is computationally more efficient than using the entire feature space ($d$).NIIL outperforms all or most of the baselines at k = 3, which can be further improved by increasing the partition size to $\frac{d}{2}$. The proposed method is evaluated on tabular data sets with categorical-only, numerical-only, and mixed data types. The sample size of these data sets ranges from 155 to 2109. Most tabular data in practice have limited sample sizes (<1000<1000) with heterogeneous data types. The imputation of data with limited samples poses an additional challenge. In contrast, deep learning studies selectively use tabular data sets with large sample sizes (>10000>10000) only [35, 8]. Therefore, our results are presented using the most practical tabular data sets, missing types, and rates to ensure generalizability.
@@ 5 Discussions
The paper proposes incremental learning for tabular data to exclude the requirement of imputing missing values. The findings of the paper can be summarized as follows. First, the proposed NIIL method outperforms all baseline methods with and without imputation regardless of the missing value types. The exclusion of missing value imputation steps makes the proposed model insensitive to missing value types, unlike other imputation-based approaches. Second, deep imputation methods may not be optimally designed to impute all missing value types, including the MNAR-type. The proposed NIIL method is a superior choice when the missing value type is unknown or complex. Third, the attention mask enables the classification of missing values without imputation. However, the performance of the AM-FTT model shows that the attention mask itself or the FTT model does not improve the classification performance. Incremental learning plays an important role along with AM-FTT to enable the proposed NIIL method to outperform the baselines. Fourth, the NIIL method offers the best performance and computational time trade-off with better resilience to missing value rates than baselines.

### v5

@@ 4 Results
All results are obtained using an Ubuntu 22.04 machine utilizing an Intel(R) Xeon(R) W-2265 CPU (24 logical cores) running at 3.70GHz, 64GB of RAM, and a Quadro RTX A4000 GPU with 16GB of video memory. The performance of the proposed IFIAL method is compared with the baseline methods using 1) the average rank based on the AUC scores, 2) the win matrix for pairwise comparisons of the methods, and 3) the robustness of the model to higher missing rates. C presents the average AUC scores for individual data sets with the MCAR type in Table 7 and the MNAR type in Table 8.
@@
Table 2: Average performance rank of the proposed (IFIAL) and baseline methods across 13 diverse MCAR-type data sets with five missing value rates.
Method Imputation 10% 20% 30% 40% 50% Avg. Rank Overall Rank
Median-GBT Yes 9.0(2.62) 7.0(4.17) 7.9(5.33) 7.2(4.46) 7.1(5.03) 7.7(0.74) 9
Median-FTT Yes 7.6(4.82) 7.9(4.42) 8.2(4.74) 7.8(5.41) 8.8(4.40) 8.1(0.42) 10
MICE-GBT Yes 7.4(3.54) 7.6(4.84) 7.5(4.47) 8.0(4.24) 7.0(4.41) 7.5(0.33) 8
MICE-FTT Yes 6.8(2.20) 6.6(2.37) 5.6(3.69) 6.4(3.05) 6.0(3.37) 6.3(0.43) 4
GAIN-GBT Yes 8.4(5.55) 10.6(3.44) 6.8(4.02) 8.4(5.13) 7.8(4.55) 8.4(1.25) 13
GAIN-FTT Yes 6.7(4.15) 6.0(4.28) 6.0(4.86) 6.9(4.67) 6.6(2.64) 6.4(0.36) 5
Diffputer-GBT Yes 9.5(3.96) 8.1(4.52) 8.5(4.90) 7.0(4.93) 7.4(4.90) 8.1(0.88) 11
Diffputer-FTT Yes 7.4(2.84) 8.6(2.06) 9.2(2.45) 9.1(1.88) 8.4(2.59) 8.5(0.66) 14
MIA-Xgboost No 8.4(4.51) 7.6(5.18) 7.8(3.96) 6.4(5.94) 6.4(5.90) 7.3(0.80) 7
MIA-LightGBM No 10.0(2.83) 9.0(2.83) 8.5(4.95) 9.0(5.66) 10.0(4.24) 9.3(0.60) 15
AM-FTT No 7.1(2.67) 8.3(3.90) 8.7(2.50) 7.9(3.18) 9.6(4.50) 8.3(0.82) 12
IFIAL (K =2) No 4.9(4.29) 4.8(2.43) 5.5(2.51) 5.4(2.83) 6.2(2.71) 5.3(0.53) 2
IFIAL (K =3) No 5.5(5.07) 6.2(3.65) 5.9(3.00) 6.9(3.56) 6.4(3.34) 6.2(0.47) 3
IFIAL (K =4) No 6.8(3.96) 6.9(4.58) 7.0(4.04) 7.4(3.62) 6.5(3.74) 6.9(0.29) 6
IFIAL (K =d/2) No 4.0(3.66) 4.1(3.60) 4.6(3.38) 4.5(2.62) 4.8(3.49) 4.4(0.29) 1
@@
Table 3: Average performance rank of the proposed (IFIAL) and baseline models across 13 diverse MNAR-type data sets with five missing value rates.
Method Imputation 10% 20% 30% 40% 50% Avg. Rank Overall Rank
Median-GBT Yes 7.1(3.73) 5.9(3.93) 6.7(4.74) 5.5(4.25) 5.1(4.86) 6.1(0.74) 7
Median-FTT Yes 6.2(3.94) 6.4(4.03) 6.4(4.20) 6.3(4.60) 6.5(4.02) 6.4(0.10) 10
MICE-GBT Yes 6.5(3.46) 7.1(4.52) 7.5(3.82) 7.6(4.00) 6.4(4.21) 7.0(0.51) 11
MICE-FTT Yes 5.4(3.62) 5.4(3.50) 4.6(3.32) 5.0(3.35) 4.6(3.43) 5.0(0.39) 3
GAIN-GBT Yes 8.4(5.55) 10.8(3.63) 7.2(4.82) 8.8(5.07) 7.6(4.56) 8.6(1.25) 14
GAIN-FTT Yes 4.9(3.62) 4.3(3.28) 4.7(3.91) 5.2(3.77) 5.0(3.35) 4.8(0.30) 2
Diffputer-GBT Yes 8.8(4.33) 7.8(4.20) 8.5(4.34) 6.8(4.40) 6.8(4.68) 7.7(0.84) 12
Diffputer-FTT Yes 5.6(3.60) 6.5(3.43) 6.5(3.72) 6.6(3.60) 5.9(3.61) 6.2(0.39) 8
MIA-Xgboost No 8.6(4.83) 8.0(4.80) 8.6(3.36) 7.2(5.81) 6.8(5.85) 7.8(0.73) 13
MIA-LightGBM No 11.5(0.71) 10.0(1.41) 10.5(2.12) 11.0(2.83) 12.0(1.41) 11.0(0.71) 15
AM-FTT No 5.9(3.33) 6.7(3.54) 6.3(2.92) 6.0(3.08) 6.9(3.92) 6.4(0.38) 9
IFIAL (K =2) No 5.4(2.61) 4.8(2.68) 5.4(2.88) 7.0(0.82) 7.0(0.82) 5.9(0.91) 5
IFIAL (K =3) No 3.8(1.79) 6.0(2.00) 4.4(2.30) 5.2(2.59) 6.0(1.87) 5.1(0.87) 4
IFIAL (K =4) No 5.8(3.56) 5.2(2.86) 6.2(2.59) 5.8(2.68) 7.2(2.95) 6.0(0.66) 6
IFIAL (K =d/2) No 3.3(2.66) 2.0(1.55) 2.8(2.64) 3.7(2.66) 5.5(2.43) 3.5(1.16) 1
@@
Table 4: Average AUC scores for four data sets with natural missing values: (Hepatitis (ID: 55), Cylinder-bands (ID: 6332), Kick (ID: 41162), Okcupid-stem (ID: 41440)) with rates of 48.8%, 48.7%, 30.2%, and 20.6%, respectively.
ID Median- Median- MICE- MICE- GAIN- GAIN- Diffputer- Diffputer- MIA- MIA- AM- IFIAL
@@ 4.1 Average performance rank
Table 2 compares the proposed IFIAL method with 11 baselines for MCAR-type data. In single-value imputation, such as median value imputation, the GBT classifier performs better than FTT, regardless of missing rates. The MICE (multiple imputation) method outperforms median-value imputation. The GAN-based imputation approach (GAIN) using the FTT classifier produces better performance than MICE at low levels of missingness (10% and 20%), but its performance deteriorates relative to MICE as the missing rate increases. In comparison, the diffusion-based imputation approach performs worse overall, with both Diffputer variants consistently appearing among the lower-performing methods across all missing rates. Among the methods without standalone imputations, the MIA-based model (MIA-XGBoost) outperforms AM-FTT, confirming the known superiority of traditional ML over DL. The average performance rank shows that the proposed IFIAL (K = d/2) outperforms all 11 baselines. IFIAL with other feature partition sizes, including IFIAL (K = 2) and IFIAL (K = 3), also rank among the top three. Even at low missing rates (10%–30%), IFIAL (K = d/2) maintains a strong performance rank, while at higher rates it shows clear superiority. These results confirm the robustness and consistent performance of the proposed IFIAL method for MCAR-type data. Table 3 reports the mean classification performance ranks for models applied to MNAR-type data. Imputation-based approaches that already perform well on MCAR data (GAIN-FTT and MICE-FTT) also achieve strong results on MNAR data, where they rank 2nd and 3rd, respectively. In contrast, the imputation-free MIA models (MIA-Xgboost and MIA-LightGBM) perform worse on MNAR data than on MCAR data. This observation suggests that while MIA-based models may benefit from informative missingness (MNAR) in large-sample size scenarios [perez2022_imputationbenchmarking], their performance can degrade when applied to relatively small data [le2025_imputation]. However, for MNAR-type data, Median-GBT consistently outperforms MIA-based methods across all missing rates, unlike in MCAR-type scenarios. In contrast, the proposed IFIAL with several partition sizes (K = 2, 3) ranks among the top five methods, while the partition (K = d / 2) produces the best performance, as it does for the MCAR-type data. These results suggest that the proposed IFIAL method is not substantially affected by missing value types, unlike other imputation-based approaches. Table 4 presents the results for the four data sets with naturally occurring missing values. In this setting, the proposed IFIAL method (with K = d/2) attains the best average performance rank (2.3 (0.02)), outperforming the second-best method, median imputed GBT classification (4.0 (0.04)). Interestingly, MIA-Xgboost shows the best on a data set (ID 6332) with limited samples but a relatively large feature space. However, for data sets with large sample sizes, differences in model performance are insignificant. The availability of a large number of samples provides sufficient information and eliminates the challenge of learning data sets with small sample sizes.
@@
Figure 3: Win matrix. Values are the fraction of experimental scenarios in which the row methods outperform the methods in the columns in terms of AUC scores.
4.2 Win matrix, computational time, and robustness
We count the experimental scenarios in which one method wins over another. There are 65 experimental scenarios for each missing value type using 13 data sets and five missing rates. Figure 3 (a) for the MCAR-type data shows that the proposed IFIAL method is superior to the baselines for more than 50% of the experimental scenarios. In the most favorable setting, IFIAL surpasses Diffputer-FTT in 89% of scenarios. Figure 3 (b), which presents the results for MNAR-type data, also shows that IFIAL surpasses all baseline methods in more than 50% of experimental scenarios. Furthermore, IFIAL outperforms Diffputer-GBT and Median-GBT in 83% of the scenarios. IFIAL also outperforms MIA-Xgboost in 76% of scenarios, which is 56% for MCAR-type data. These comparisons indicate that deep-imputation methods are not optimized for MNAR-type data.
@@ 4.3 Computational time and robustness to missing rates
Table 5 presents the run time for the imputation and classification steps of the baseline methods. As expected, methods faster than IFIAL include GAIN-GBT, Diffputer-GBT, AM-FTT, MIA-Xgboost, and MIA-LightGBM, mainly due to the fast GBT classifier model. However, the classification performance rank of these methods is worse than that of IFIAL. One computational benefit of IFIAL is that it saves the time required for imputation. Figure 4 shows the effect of varying the missing rates on the performance of five top performing methods. The baseline AUC values used in the robustness analysis are derived from classifications performed on full data sets with 0% missing values. The AUC score decreases relative to the reference AUC as the missing rate increases. Figure 4(a) for MCAR-type data shows better robustness of the proposed method when the missing rate is greater than 30%. However, Figure 4(b) for MNAR-type data demonstrates that the proposed IFIAL method preserves its robustness even as the missing rate increases, clearly outperforming all other competitive baselines.
@@
MICE - GBT 417 2 419 8 11
MICE - FTT 417 190 607 4 3
GAIN - GBT 51 2 54 13 14
GAIN - FTT 51 190 241 5 2
Diffputer - GBT 8 1 9 11 12
Diffputer - FTT 8 190 198 14 8
AM-FTT N/A 21 21 12 9
MIA - Xgboost N/A 1 1 7 13
MIA - LightGBM N/A 1 1 15 15
IFIAL (k = $\frac{d}{2}$) N/A 163 163 1 1
@@ 4.4 Ablation studies
In an ablation study, we vary the feature partition size (k) for three representative data sets: Kc2 (ID 1063), Diabetes (ID 37), and Dresses sales (ID 23381). Figure 5 shows the k values that yield the best AUC scores for individual data sets. In general, incremental learning of features is superior to learning with all features at once (that is, without incremental learning, k = d). The best average AUCs are obtained for the data sets Kc2 $(d=21)$ at k=10, Diabetes $(d=8)$ at k=6, and Dresses sales $(d=12)$ at k = 6. Therefore, the value of k can be chosen as half of the feature dimension for the best classification performance. This observation aligns with the computational cost analysis presented in Figure 6 in  B, where $k<\frac{d}{2}$ is computationally more efficient than using the entire feature space ($d$). In general, IFIAL outperforms all or most of the baselines at k = 3, which can be further improved by increasing the size of the partition to $\frac{d}{2}$.
@@ 5 Discussions
The paper proposes incremental learning for tabular data to eliminate the requirement of initializing, imputing, and completing missing values. The findings of the paper can be summarized as follows. First, the proposed IFIAL method outperforms all baseline methods with and without data imputation, regardless of the missing value types. Second, state-of-the-art imputation methods are sensitive to missing value types, especially when missing values are not at random (MNAR). The proposed IFIAL method entirely avoids the imputation of missing values and, therefore, is a superior choice when the missing value type is unknown or complex. Third, while retrofitting a pair of attention masks enables imputation-free learning of tabular data representations, incremental learning of feature partitions is essential for overall efficiency and performance. In comparison, the performance of AM-FTT suggests that attention masks and the FTT model are insufficient to achieve optimal performance. Fourth, the IFIAL method offers the best performance and computational time trade-off with better resilience to missing value rates than baselines. In general, our proposed method renders existing missing value imputation methods unnecessary from both computational efficiency and performance standpoints.